### v1

# Hyperbolic Fuzzy $C$-Means with Adaptive Weight-based Filtering for Clustering in Non-Euclidean Spaces

Abstract
Clustering algorithms play a pivotal role in unsupervised learning by identifying and grouping similar objects based on shared characteristics. While traditional clustering techniques, such as hard and fuzzy center-based clustering, have been widely used, they struggle with complex, high-dimensional, and non-Euclidean datasets. In particular, the Fuzzy $C$-Means (FCM) algorithm, despite its efficiency and popularity, exhibits notable limitations in non-Euclidean spaces. Euclidean spaces assume linear separability and uniform distance scaling, limiting their effectiveness in capturing complex, hierarchical, or non-Euclidean structures in fuzzy clustering. To overcome these challenges, we introduce Filtration-based Hyperbolic Fuzzy $C$-Means (HypeFCM), a novel clustering algorithm tailored for better representation of data relationships in non-Euclidean spaces. HypeFCM integrates the principles of fuzzy clustering with hyperbolic geometry and employs a weight-based filtering mechanism to improve performance. The algorithm initializes weights using a Dirichlet distribution and iteratively refines cluster centroids and membership assignments based on a hyperbolic metric in the Poincaré Disc model. Extensive experimental evaluations demonstrate that HypeFCM significantly outperforms conventional fuzzy clustering methods in non-Euclidean settings, underscoring its robustness and effectiveness.
Index Terms:
Clustering, Hyperbolic Geometry, Fuzzy $C$-means (FCM), Poincaré Disc.
1 Introduction.
Clustering is an unsupervised learning technique used to group objects based on similarities, organizing similar items into clusters to reveal underlying patterns or structures. Clustering is a powerful tool for analyzing complex datasets by identifying meaningful groupings without predefined labels. A wide range of clustering algorithms has been proposed, broadly distinguished into hard and soft clustering methods. While hard clustering assigns each data point to a single cluster, soft clustering allows overlaps. Hard clustering methods include center-based, hierarchical-based, distribution-based, and density-based techniques. Center-based methods, such as $k$-means [1], $k$-medoids [2], $k$-harmonic means [3], spectral clustering [4], kernel $k$-means [5] measure similarity by proximity to cluster centers. Hierarchical-based methods, including hierarchical clustering [6], agglomerative clustering [7] assume stronger similarities between closer data points. Distribution-based methods, such as Expectation Maximization (EM) for Gaussian Mixture Models [8], robust EM for Gaussian Mixture Models [9], cluster the data based on probability distributions. Density-based methods, like DBSCAN [10], HDBSCAN [11], Mean-Shift [12] identify clusters by analyzing data density in feature space.
Soft clustering allows data points to be associated with multiple clusters with varying degrees of membership. Cluster labels are based on the highest membership value. Notable techniques include possibilistic clustering [13] and fuzzy clustering [14], with fuzzy $c$-means (FCM) [15], fuzzy density peaks clustering [16], centroid auto-fused hierarchical FCM [17], robust FCM [18].
FCM is widely used for its efficiency and simplicity, yet it struggles with complex, high-dimensional, and non-Euclidean datasets. To mitigate these limitations, several variants have been introduced, incorporating improved objective functions and constraints, such as adaptive FCM [19], generalized FCM [20], fuzzy weighted $c$-means [21], and generalized FCM with improved fuzzy partitioning [22]. Kernel-based approaches like kernel FCM (KFCM) [5] and constrained models, including agglomerative fuzzy $k$-means (AFKM) [23], robust self-sparse fuzzy clustering (RSSFCA) [18], robust and sparse fuzzy $k$-means (RSFKM) [24], possibilistic FCM (PFCM) [25], and principal component analysis-embedded FCM (P_SFCM) [26] as well as hyperbolic extensions such as hyperbolic smoothing-based fuzzy clustering (HSFC) [27] and Integration of hyperbolic tangent and Gaussian kernels for FCM (HGFCM) [28], have also been explored. However, despite these modifications, most of these methods remain fundamentally limited in non-Euclidean spaces, as they can still partially rely on Euclidean assumptions that fail to capture the inherent geometric complexity and hierarchical structure of such data. Consequently, these approaches often exhibit suboptimal clustering performance, reinforcing the need for a more robust, geometry-aware solution.
To overcome such challenges, we propose Filtration-based Hyperbolic Fuzzy $\mathbf{C}$-Means or HypeFCM, an efficient fuzzy clustering algorithm for non-Euclidean spaces. Our approach integrates fuzzy clustering principles with hyperbolic geometry and employs a weight-based filtering mechanism to improve cluster assignments. The algorithm initializes weights using a Dirichlet distribution, iteratively updates centroids and membership weights, and computes distances using the metric of the Poincaré Disc model. The Poincaré Disc metric naturally preserves hierarchical structures due to the exponential distance scaling and the negative curvature of hyperbolic space. In contrast to Euclidean space, where distances grow linearly, hyperbolic distances expand exponentially as points move away from the origin. This property allows hyperbolic space to efficiently embed tree-like and hierarchical structures, as the distances between levels of a hierarchy are naturally stretched, ensuring clear separability between different clusters. Moreover, the Poincaré Disc model maintains conformal mapping, meaning local angles and relative positioning are preserved, which helps retain structural integrity when clustering complex, high-dimensional data. These characteristics make the Poincaré Disc metric particularly well-suited for capturing latent relationships and hierarchical patterns in non-Euclidean datasets where traditional Euclidean methods fall short. A filtration step refines connections by focusing on relevant associations, followed by weight updates until convergence. The algorithm culminates in a defuzzification process, where definitive cluster assignments are determined by optimizing the final weight matrix, effectively mapping the continuous membership values to discrete cluster labels. Our approach leverages hyperbolic space properties to improve clustering performance in complex, high-dimensional datasets.
Contributions. The key contributions are:
  1. 1.
    Our proposed algorithm HypeFCM incorporates the metric of the Poincaré Disc model by embedding the datasets inside the Poincaré Disc for more efficient performance in non-Euclidean spaces.
  2. 2.
    Our method introduces a selective filtration process that prunes less significant relationships, optimizes computational efficiency by focusing on the most relevant geometric connections, and improves cluster definition through a targeted weight refinement strategy.
  3. 3.
    We present a thorough convergence analysis and discuss the computational complexity analysis of the proposed method. Other necessary proofs and derivations are provided in the supplementary document.
The paper is structured as follows: Section 2 introduces the mathematical preliminaries, Section 3 reviews related methods, and Section 4 presents the proposed HypeFCM algorithm in detail. Section 5 analyzes its computational complexity, while Section 6 discusses convergence properties. Section 7 showcases experimental results on real and synthetic datasets, Section 8 contains detailed discussions of the proposed method, and finally, Section 9 concludes the paper.
2 Preliminaries.
We present the fundamental mathematical preliminaries of hyperbolic geometry required for our work. For a more in-depth exploration, refer to [29], [30].
2.1 Smooth Manifolds.
2.1.1 Manifold.
A topological space $\mathcal{M}$ is locally Euclidean of dimension $d$, if for every $\sigma\in\mathcal{M}$, there exists an open neighborhood $(X_{\sigma})$ and a map $\psi_{\sigma}:X_{\sigma}\rightarrow\psi_{\sigma}(X_{\sigma})$, where $\psi_{\sigma}(X_{\sigma})\subset\mathbb{R}^{d}$ is a homeomorphism. An $d$-dimensional Manifold is Hausdorff, $2^{nd}$ countable and locally Euclidean of dimension $d$.
  • Each $(X_{\sigma},\psi_{\sigma})$ is called a chart.
  • The family $\{(X_{\sigma},\psi_{\sigma})\}_{\sigma=1}^{k}$ is an atlas.
2.1.2 Smooth Manifold.
A smooth or $\mathbb{C}^{\infty}$ manifold is a topological manifold which comes with a family $(X_{\sigma},\psi_{\sigma})$ with the sets $X_{\sigma}\subset\mathcal{M}$ and the homeomorphisms $\psi_{\sigma}:X_{\sigma}\rightarrow\mathbb{R}^{d}$ such that $\mathcal{M}=\cup_{\sigma}X_{\sigma}$ given $(X_{\sigma}\cap X_{\beta})$ is non-null and the map $(\psi_{\beta}\circ\psi_{\sigma}^{-1})$ is smooth.
2.2 Riemannian Manifold
A Riemannian manifold $(\mathcal{M},g)$, consists of a smooth $d$-dimensional manifold $\mathcal{M}$ equipped with a Riemannian metric $g$ on $\mathcal{M}$.
For each point $p\in\mathcal{M}$, $g$ assigns a bilinear form $g_{p}:T_{p}(\mathcal{M})\times T_{p}(\mathcal{M})\rightarrow\mathbb{R}$, where $T_{p}(\mathcal{M})$ is the tangent space at $p$, such that:
  • $g_{p}$ is symmetric: $g_{p}(A,B)=g_{p}(B,A)$ for all $A,B\in T_{p}(\mathcal{M})$.
  • $g_{p}$ is positive-definite: $g_{p}(A,A)>0$ for all non-zero $A\in T_{p}(\mathcal{M})$.
If $(x^{1},x^{2},\ldots,x^{d}):U\subset\mathcal{M}\rightarrow\mathbb{R}^{d}$ are smooth local coordinates on $\mathcal{M}$, the vectors $\left\{\left(\frac{\partial}{\partial x^{1}}\right)_{p},\left(\frac{\partial}{\partial x^{2}}\right)_{p},\ldots,\left(\frac{\partial}{\partial x^{d}}\right)_{p}\right\}$ form a basis of the vector space $T_{p}(\mathcal{M})$ for any $p\in U$.
2.2.1 Length and Distance
[…]
Let $(\mathcal{M},g)$ be a connected Riemannian Manifold. Given two points $\mathbf{a,b}\in\mathcal{M}$, the distance between $\mathbf{a,b}$ is defined as: $d_{g}(\mathbf{a,b})=inf\{L(\gamma):\gamma$ is a piecewise smooth curve with $\gamma(0)=\mathbf{a},\gamma(1)=\mathbf{b}\}$. This distance is called the geodesic distance.
Let $\varsigma(\mathcal{M})$ denote the space of all $\mathbb{C}^{\infty}$ vector fields on $\mathcal{M}$. Let $\nabla:\varsigma(\mathcal{M})\times\varsigma(\mathcal{M})\rightarrow\varsigma(\mathcal{M})$ be a bi-linear map, therefore the following must hold,
  • $\nabla_{(lX_{1}+l^{\prime}X_{2})}X_{3}=l\nabla_{X_{1}}X_{3}+l^{\prime}\nabla_{X_{2}}X_{3},\forall X_{1},X_{2},X_{3}\in\varsigma(\mathcal{M})$ and $l,l^{\prime}\in\mathbb{C}^{\infty}(\mathcal{M}).$
  • $\nabla_{X_{1}}lX_{2}=X_{1}(l)X_{2}+l\nabla_{X_{1}}X_{2},\forall l\in\mathbb{C}^{\infty}(\mathcal{M}).$
$\nabla$ is said to preserve the metric and torsion-free if it satisfies the following,
  • $X_{1}(g(X_{2},X_{3}))=g(\nabla_{X_{1}}X_{2},X_{3})+g(X_{2},\nabla_{X_{1}}X_{3})$.
  • $\nabla_{X_{1}}X_{2}-\nabla_{X_{2}}X_{1}=[X_{1},X_{2}]$.
Then, $\nabla$ on $(\mathcal{M},g)$ is called the Levi-Civita connection.
2.2.2 Curvature Operator
The Riemannian curvature $\mathcal{R}$ on $\mathcal{M}$ is a trilinear map defined by: $\mathcal{R}((X_{1},X_{2}),X_{3})=\nabla_{X_{1}}\nabla_{X_{2}}X_{3}-\nabla_{X_{2}}\nabla_{X_{1}}X_{3}-\nabla_{[X_{1},X_{2}]}X_{3}\quad\forall X_{1},X_{2},X_{3}\in\varsigma(\mathcal{M}).$
2.2.3 Sectional Curvature
Let $\pi=(p,\mathscr{X})$, where $p\in\mathcal{M}$ and $\mathscr{X}\subset T_{p}(\mathcal{M})$ be a two-dimensional subspace spanned by ${A_{1},A_{2}}$. The sectional curvature of $\mathcal{M}$ is defined as,
$\mathbf{k}{(\pi)}=\frac{\langle\mathcal{R}(A_{1},A_{2})A_{2},A_{1}\rangle}{\langle A_{1},A_{1}\rangle\langle A_{2},A_{2}\rangle-\langle A_{1},A_{2}\rangle}.$
2.3 Hyperbolic space
A Hyperbolic space, represented as $\mathbb{H}^{n}$, is a non-Euclidean space of dimension $n$ which is characterized as a simply-connected Riemannian manifold with a constant negative sectional curvature $-1$. The Killing-Hopf Theorem [31] affirms that any two such Riemannian manifolds are isometrically equivalent. We will briefly discuss four of them here.
  1. 1.
    Poincaré Half Space Model: This is the upper half plane $\mathbb{H}^{n}:=\left\{\left(x_{1},x_{2},\ldots,x_{n}\right)\mid x_{n}>0,x_{i}\in\right.$ $\mathbb{R},\forall i\in\{1,2,\ldots,n\}\}$. For $\mathbf{X}:=\left(x_{1},x_{2},\ldots,x_{n}\right)$ and $\mathbf{Y}:=\left(y_{1},y_{2},\ldots,y_{n}\right)$. $\mathbf{\tilde{X}}:=\left(x_{1},x_{2},\ldots,x_{n-1},-x_{n}\right)$, the reflection of $\mathbf{X}$ with respect to the plane $x_{n}=0$ then, the metric on this space is defined as follows:
    $d\left(\mathbf{X},\mathbf{Y}\right)$ $:=2\sinh^{-1}\frac{\left\|\mathbf{Y}-\mathbf{X}\right\|}{2\sqrt{x_{n}y_{n}}}$ (1)
    $=2\log\frac{\left\|\mathbf{Y}-\mathbf{X}\right\|+\left\|\mathbf{Y}-\mathbf{\tilde{X}}\right\|}{2\sqrt{x_{n}y_{n}}}.$
  2. 2.
    Poincaré Disc Model: This is a model of hyperbolic space in which all points are inside the unit disc in $\mathbb{R}^{n}$, and geodesics are either diameters or the circular arcs. The metric between two points $\mathbf{X}$ and $\mathbf{Y}\left(\left\|\mathbf{X}\right\|,\left\|\mathbf{Y}\right\|<1\right)$ is defined as,
    $d\left(\mathbf{X,Y}\right):=\cosh^{-1}\left(1+\frac{2\left\|\mathbf{X}-\mathbf{Y}\right\|^{2}}{\left(1-\left\|\mathbf{X}\right\|^{2}\right)\left(1-\left\|\mathbf{Y}\right\|^{2}\right)}\right).$ (2)
  3. 3.
    Beltrami-Klein Model: This is a model hyperbolic space in which points are represented by the points in the open unit disc in $\mathbb{R}^{n}$ and lines are represented by chords, a special type of straight lines with ideal endpoints a, b on the unit sphere. For $\mathbf{x,y}\in\mathbb{D}^{n}$, the open unit ball, the distance in this model is given as
    $d(\mathbf{x},\mathbf{y}):=\frac{1}{2}\log\left(\frac{\|\mathbf{a}\mathbf{y}\|\|\mathbf{x}\mathbf{b}\|}{\|\mathbf{a}\mathbf{x}\|\|\mathbf{y}\mathbf{b}\|}\right),$ (3)
    where $\mathbf{a,b}$ are ideal points on the boundary of the sphere of the line $(\mathbf{x,y})$ with $\|\mathbf{a}\mathbf{y}\|>\|\mathbf{a}\mathbf{x}\|$ and $\|\mathbf{x}\mathbf{b}\|>\|\mathbf{b}\mathbf{y}\|$.
  4. 4.
    Hyperboloid Model: This is a model of hyperbolic space also known as the Minkowski Model, which is the forward sheet $S^{+}$of the two hyperbolic sheets embedded in the $(n+1)$ dimensional Minkowski Space. The hyperbolic distance between two points $\mathbf{X}=\left(x_{0},x_{1},\ldots,x_{n}\right)$ and $\mathbf{Y}=\left(y_{0},y_{1},\ldots,y_{n}\right)$ is given as,
    $d(\mathbf{X,Y}):=\cosh^{-1}(-M(\mathbf{X,Y})).$ (4)
    where,
    $M\left(\left(x_{0},x_{1},\ldots,x_{n}\right),\left(y_{0},y_{1},\ldots,y_{n}\right)\right)$
    $\qquad\qquad=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
    Here, $M$ is simply the Minkowski Dot Product.
2.4 Gyrovector spaces.
The framework of gyrovector spaces establishes an elegant non-associative algebraic structure that naturally captures and formalizes the properties of hyperbolic geometry analogous to the vector spaces provide the algebraic setting for Euclidean geometry [32]. We denote $\mathbb{D}_{\alpha}^{p}:=\left\{\mathbf{v}\in\mathbb{R}^{p}\mid\alpha\|\mathbf{v}\|^{2}<1\right\}$ taking $\alpha\geq 0$. If $\alpha=0$, then $\mathbb{D}_{\alpha}^{p}=\mathbb{R}^{p}$; if $\alpha>0$, then $\mathbb{D}_{\alpha}^{p}$ is the open ball of radius $1/\sqrt{\alpha}$. If $\alpha=1$ then we recover the unit ball $\mathbb{D}^{p}$.
The operations in Poincaré Disc Model are well-defined, computationally efficient, as most of the geometric operations have simple closed-form expressions in Poincaré Disc Model, and it also preserves the angle. In this context, we will provide a brief discussion on Möbius gyrovector addition and Möbius scalar multiplication on the Poincaré Disc model. Due to the invariance of geometric properties under isometric mappings between the hyperbolic spaces, the fundamental additive and multiplicative algebraic structures can be isomorphically translated across different models of hyperbolic geometry while preserving their essential characteristics [33]. We have established the isometric equivalence between the Poincaré Disc model and the hyperboloid model by explicitly constructing a mapping between them. Let $\mathbb{D}^{n}=\{x\in\mathbb{R}^{n}:\|x\|<1\}$ be the Poincaré Disc model, and let the hyperboloid model be
$\mathbb{H}^{n}=\{x\in\mathbb{R}^{n+1}:-x_{0}^{2}+x_{1}^{2}+\cdots+x_{n}^{2}=-1,\ x_{0}>0\}.$
Then the mapping:
$\phi:\mathbb{D}^{n}\to\mathbb{H}^{n},\quad x\mapsto\left(\frac{1+\|x\|^{2}}{1-\|x\|^{2}},\ \frac{2x}{1-\|x\|^{2}}\right)$
is an isometry. It preserves the hyperbolic distance and induces the same Riemannian metric. Hence, we will be using the Poincaré Disc model throughout.
Möbius addition. The Möbius addition of $\mathbf{v}$ and $\mathbf{w}$ in $\mathbb{D}_{\alpha}^{p}$ is defined as :
$\mathbf{v}\oplus_{\alpha}\mathbf{w}:=\frac{\left(1+2\alpha\langle\mathbf{v},\mathbf{w}\rangle+\alpha\|\mathbf{w}\|^{2}\right)\mathbf{v}+\left(1-\alpha\|\mathbf{v}\|^{2}\right)\mathbf{w}}{1+2\alpha\langle\mathbf{v},\mathbf{w}\rangle+\alpha^{2}\|\mathbf{v}\|^{2}\|\mathbf{w}\|^{2}}.$ (5)
In particular, when $\alpha=0$, this conforms with the Euclidean addition of two vectors in $\mathbb{R}^{p}$. However, it satisfies $\mathbf{v}\oplus_{\alpha}\mathbf{0}=\mathbf{0}\oplus_{\alpha}\mathbf{v}=\mathbf{v}$. Moreover, for any $\mathbf{v},\mathbf{w}\in\mathbb{D}_{\alpha}^{p}$, we have $(-\mathbf{v})\oplus_{\alpha}\mathbf{v}=\mathbf{v}\oplus_{\alpha}(-\mathbf{v})=\mathbf{0}$ and $(-\mathbf{v})\oplus_{\alpha}\left(\mathbf{v}\oplus_{\alpha}\mathbf{w}\right)=\mathbf{w}$.
Möbius scalar multiplication. For $\alpha>0$, the Möbius scalar multiplication of $\mathbf{v}\in\mathbb{D}_{\alpha}^{p}\backslash\{\boldsymbol{0}\}$ by a real number $\lambda\in\mathbb{R}$ is defined as,
$\lambda\otimes_{\alpha}\mathbf{v}:=(1/\sqrt{\alpha})\tanh\left(\lambda\tanh^{-1}(\sqrt{\alpha}\|\mathbf{v}\|)\right)\frac{\mathbf{v}}{\|\mathbf{v}\|}.$ (6)
and $\lambda\otimes_{\alpha}\mathbf{0}:=\mathbf{0}$. As the parameter $\alpha$ approaches zero, the expression reverts to conventional Euclidean scalar multiplication: $\lim_{\alpha\rightarrow 0}\lambda\otimes_{\alpha}\mathbf{v}=\lambda\mathbf{v}$.
The Hyperbolic Distance function on $\left(\mathbb{D}_{\alpha}^{p},g^{\alpha}\right)$ is given by,
$d_{H}(\mathbf{v},\mathbf{w})=(2/\sqrt{\alpha})\tanh^{-1}\left(\sqrt{\alpha}\left\|-\mathbf{v}\oplus_{\alpha}\mathbf{w}\right\|\right).$ (7)
Riemannian Log-Exp Map. Given two points $x,y\in\mathbb{D}_{\alpha}^{p}$, the logarithmic map at $x$ applied to $y$ is denoted as:
$\log_{\mathbf{x}}(\mathbf{y})=\frac{2}{\lambda_{\mathbf{x}}}\cdot\frac{\tanh^{-1}\left(\sqrt{\alpha}\cdot\|-\mathbf{x}\oplus_{\alpha}\mathbf{y}\|\right)}{\|-\mathbf{x}\oplus_{\alpha}\mathbf{y}\|}\cdot\left(-\mathbf{x}\oplus_{\alpha}\mathbf{y}\right),$ (8)
where ${\lambda_{\mathbf{x}}}=\frac{2}{1-\lambda{\|x\|}^{2}}$ and $\bigoplus_{\alpha}$ represents the Möbius gyrovector addition as Equation 5, $n$ is the number of data points, $m$ is the fuzziness parameter.
Given $x\in\mathbb{D}_{\alpha}^{p}$, $v\in T_{X}(\mathbb{D}_{\alpha}^{p})$, the exponential map is defined as,
$\exp_{\mathbf{x}}(\mathbf{v})=\mathbf{x}\oplus_{\alpha}\left(\tanh\left(\frac{\lambda_{\mathbf{x}}\cdot\sqrt{\alpha}\cdot\|\mathbf{v}\|}{2}\right)\cdot\frac{\mathbf{v}}{\sqrt{\alpha}\cdot\|\mathbf{v}\|}\right).$ (9)
As the parameter $\alpha$ approaches to 00,
$\lim_{\alpha\rightarrow 0}\exp_{\mathbf{x}}(\mathbf{v})=\mathbf{(x+v)},$ (10)
$\lim_{\alpha\rightarrow 0}\log_{\mathbf{x}}(\mathbf{y})=\mathbf{(y-x)},$ (11)
which are the exponential and logarithmic maps in Euclidean space.
3 related works.
3.1 Fuzzy $C$-means.
The most well-known method for fuzzy clustering is the FCM clustering algorithm[15]. Let $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots\mathbf{x}_{n}\}$ be a sample of $n$ observations in $\mathbb{R}^{p}$. Thus, $\mathbf{X}$ is a $(n\times p)$ data matrix, $\mathbf{V}$ is an $(c\times p)$ matrix that represents centroids of the $c$ clusters in $\mathbb{R}^{p}$ and $\mathbf{W}$ is the $(n\times c)$ membership matrix with elements $w_{ij}\in[0,1]$, then the following constrained weighted least square criterion, which is also the objective function for FCM is to be minimized:
$Minimize\hskip 5.0ptJ(\mathbf{X},\mathbf{W},\mathbf{C})=\sum_{i=1}^{n}\sum_{j=1}^{c}(w_{ij})^{m}\|\mathbf{x}_{i}-\mathbf{v}_{j}\|^{2},$
subject to
$\sum_{j=1}^{c}w_{ij}=1,\quad\text{for all }i\in\{1,2,\dots,n\},$
$0<\sum_{i=1}^{n}w_{ij}<n,\quad\text{for all }j\in\{1,2,\dots,c\}.$ (12)
where, $\mathbf{x}_{i}$ is the $i^{th}$ row of $\mathbf{X}$, $\mathbf{v}_{j}$ is the $j^{th}$ row of $\mathbf{V}$ in $\mathbb{R}^{p}$, the centroid of the $j^{th}$ cluster.
The parameter $m\in[2,\infty)$ in [34] is the fuzziness parameter. According to [35], the value of $m$ is usually taken as 2. The optimization problem in Equation 12 can be solved using Lagrange multipliers [36]. Using the Alternative Minimization method, the update formula for the centroid is given by :
$\mathbf{v}_{j}=\frac{\sum_{i=1}^{n}(w_{ij})^{m}\mathbf{x}_{i}}{\sum_{i=1}^{n}(w_{ij})^{m}}.$ (13)
The solution for the membership matrix $\mathbf{W}$ is given as:
$w_{ij}=\left[\sum_{k=1}^{c}\left(\frac{\|\mathbf{x}_{i}-\mathbf{v}_{j}\|^{2}}{\|\mathbf{x}_{i}-\mathbf{v}_{k}\|^{2}}\right)^{\frac{1}{m-1}}\right]^{-1}.$ (14)
3.2 HSFC
For the clustering problem of the $n$ rows of data matrix $\mathbf{X}$ in $c$ clusters, we can seek for the minimum distance between every $\mathbf{x}_{i}$ and its cluster centroid $\mathbf{v}_{j}$ :
$d_{i}^{2}=\min_{\mathbf{v}_{j}\in\mathbf{V}}\left\|\mathbf{x}_{i}-\mathbf{v}_{j}\right\|_{2}^{2}.$
Consider the hyperbolic smoothing function $\phi_{(}y,\tau)=\frac{y+\sqrt{y^{2}+\tau^{2}}}{2}$, for all $y\in\mathbb{R},\tau\geq 0$ and the function: $\theta\left(\mathbf{x}_{i},\mathbf{v}_{j},\gamma\right)=$ $\sqrt{\sum_{k=1}^{p}\left(x_{ik}-v_{jk}\right)^{2}+\gamma^{2}}$, for $\gamma>0$. Hence, the Minimization Problem of Hyperbolic Smoothing Clustering Method (HSFC) [37] is transformed into:
$\begin{array}[]{ll}\min\hskip 6.0pt\sum_{i=1}^{n}d_{i}^{2},\\ \text{subject to}\sum_{j=1}^{c}\psi_{\sigma}\left(d_{i}-\theta\left(\mathbf{x}_{i},\mathbf{v}_{j},\gamma\right),\tau\right)\geq\epsilon,\quad\forall i\in[n].\end{array}$ (15)
3.3 P_SFCM
Assume a data matrix is indicated as $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n$ is the number of data points and $p$ indicates the dimensionality of the data. $\mathbf{W}=(w_{ij})$ is the $(n\times c)$ membership matrix. Hence, the Minimization Problem of P_SFCM is given as :
$\min_{\mathbf{P},\mathbf{W},\mathbf{V}}\sum_{i=1}^{n}f_{i}\sum_{j=1}^{c}\|\mathbf{x}_{i}\mathbf{P}^{T}-\mathbf{\tilde{v}}_{j}\|_{2}^{2}(w_{ij})^{m}-\lambda Tr(\mathbf{P}\mathbf{S}_{t}\mathbf{P}^{T}),$
[…]
where, $\mathbf{P}\in\mathbb{R}^{q\times p}$ is the projection matrix, such that $\mathbf{X}\mathbf{P}^{T}\in\mathbb{R}^{n\times q},q\leq p$, sum of each row of $\mathbf{W}$ is 1, $\mathbf{\tilde{v}}_{j}$ denotes the cluster centriods, $f_{i}$ is the element of the matrix $\mathbf{F}$ s.t. $f_{i}\in\{0,1\}.$
4 HypeFCM Method and its algorithm.
In this section, we will present our proposed clustering algorithm, HypeFCM with intricate details.
Embedding into a Poincaré Disc. Our algorithm begins with embedding the dataset X(n×p) into the hyperbolic space, here the Poincaré Disc model of radius $\frac{1}{\sqrt{\alpha}}$. We obtain $\mathbf{X^{\prime}}=\{\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime},\cdots,\mathbf{x}_{n}^{\prime}\}\in\mathbb{D}_{\alpha}^{p}$, where $\mathbb{D}_{\alpha}^{p}:=\left\{\mathbf{x}\in\mathbb{R}^{p}\mid\alpha\|\mathbf{x}\|^{2}<1\right\}$.
Initializing the Weight Matrix. The weight matrix $\textbf{W}=(w_{ij})_{n\times c}$, contains the membership probabilities of the corresponding datapoint of belonging to one of the clusters. Here, the rows of W are initialized as i.i.d. samples from the Dirichlet distribution with all equal probabilities, i.e.
${\{w_{ij}\}_{j=1}^{c}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Dirichlet(\mathbf{1}_{c}/(\mathbf{1}^{T}\mathbf{1}))}\quad\forall i\in[n],$
where, $\mathbf{1}_{c}$ is the vector $(1,1,\cdots,1)\in\mathbb{R}^{c}$.
Input: Data (Xn×p), $k$ = filtration value, $c$ = number of clusters, $m$ = fuzziness parameter, $T$ = maximum number of iterations, $\varepsilon$ = tolerance value.
Output: The clusters $Y_{1},Y_{2},\cdots Y_{c}.$
1: Embed the original dataset $\mathbf{X}$ into the Poincaré Disc of radius $\frac{1}{\sqrt{\alpha}}$. Obtain $\mathbf{X^{\prime}}=\{\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime},\cdots,\mathbf{x}_{n}^{\prime}\}\in\mathbb{D}_{\alpha}^{p}.$
[…]
3: Obtain the centroids $\{\mathbf{v}^{(t)}_{j}\}_{j=1}^{c}$ using the Riemannian log-exp map:
$\bar{v}_{j}=\frac{\sum_{i=1}^{n}(w_{ij}^{(t)})^{m}\cdot\log_{\mathbf{v}_{j}^{(t-1)}}(\mathbf{x}_{i}^{\prime})}{\sum_{i=1}^{n}(w_{ij}^{(t)})^{m}}$
$\mathbf{v}_{j}^{(t)}=\exp_{\mathbf{v}_{j}^{(t-1)}}(\bar{v}_{j})$
where, $\log$ and $\exp$ denote the logarithmic and exponential maps on the Poincaré ball respectively.
4: Obtain the distance matrix, U(t) by considering distances from the centroids,
$\mathbf{U^{(t)}_{n\times c}}:=u^{(t)}_{ij}:=d_{H}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}^{(t)}_{j}))_{i\in[n],j\in[c]}.$
using the squared hyperbolic distance between data points and centroids.
5: Apply filtration based on the distances of data points from each of the centroids and the rest of the entries of U(t), are assigned to zero to obtain $\mathbf{U}^{\prime(t)}$.
6: Update the weights using:
$w^{(t+1)}_{ij}=\frac{(u^{\prime(t)}_{ij})^{-1/(m-1)}}{\sum_{j}(u^{\prime(t)}_{ij})^{-1/(m-1)}}.$
7:for $t$ from $1$ to $T$ do
       Update $\mathbf{W}^{(t)}$ by Step ($6$)
      Update centroids $\mathbf{V}^{(t)}:=\{\mathbf{v}^{(t)}_{j}\}_{j=1}^{c}$ by Step ($3$)
      if $||\mathbf{W}^{(t+1)}-\mathbf{W}^{(t)}||\leq\varepsilon$ then
            break
       end if
end for
8: Assign labels by selecting the index of the maximum membership value in each row of the finally updated weight matrix.
Algorithm 1 HypeFCM Algorithm (Filtration-based Hyperbolic Fuzzy $C$-means)
Computing the Centroids. The matrix (Vc×p), contains $c$ different centroids in each of its rows. The update expression for the centroids at each iteration is given as:
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}^{\prime}_{i})}{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}}\right).$ (17)
When limit $\alpha$ tends to 00, this update formula 17 becomes:
$\textbf{v}_{j}^{(t+1)}\xrightarrow{\alpha\to 0}\textbf{v}_{j}^{(t)}+\frac{\sum_{i}(w_{ij})^{m}(\mathbf{x}^{\prime}_{i}-\textbf{v}_{j}^{(t)})}{\sum_{i}(w_{ij})^{m}}=\frac{\sum_{i}(w_{ij})^{m}\textbf{x}^{\prime}_{i}}{\sum_{i}(w_{ij})^{m}},$
which is the update formula for the centroid in FCM in Euclidean space using the equations 10, 11.
Computing the distance matrix and applying filtration. Obtain a distance matrix $\mathbf{U}$,
$\mathbf{U^{(t)}_{n\times c}}:=u^{(t)}_{ij}:=d_{H}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}^{(t)}_{j})_{i\in[n],j\in[c]},$
where, $d_{H}$ is the hyperbolic distance mentioned in Equation 7. In each row of the matrix $\mathbf{U}$, the nearest $k$ data points from the respective centroid are kept as it is and the remaining entries are assigned to 00, that is, given the $t^{th}$ iteration of the centroids, only those data points that are the least distant from the respective centroid are kept and the rest are zero.
In essence, we define a relation $\mathbf{R}$ on the Cartesian product of the set of the centroids and the set of the datapoints, where a datapoint $i$ and a centroid $j$ are related such that,
$(i,j)\in\mathbf{R}\iff u_{ij}\leq u_{i^{\prime}j},\forall i^{\prime}\in S\subset[n]\bigg{|}n\geq|S|\geq(n-k).$
Updating the Weight Matrix and Centroids. We construct an updated distance matrix, say $\mathbf{U}^{\prime}_{n\times c}$, where the $(i,j)^{th}$ entry contains the updated distance between $\mathbf{x}_{i}^{\prime}$ and $\mathbf{v}_{j}$ after applying the filtration. The expression for this is given as:
$u^{\prime}_{ij}=\Bigl{(}d_{H}^{2}(\mathbf{x}_{i}^{\prime}-\mathbf{v}_{j})\mathds{1}\{(i,j)\in\mathbf{R}\}\Bigr{)}\quad\forall i\in[n],\forall j\in[c],$
where ${.}\mathds{1\{.\}}$ denotes the indicator function and $\mathbf{R}$ being the relation as described above. Then the expression for the updation of the weight matrix $\mathbf{W}$ is given as:
$w_{ij}^{(t+1)}=\frac{({u^{\prime}}_{ij}^{(t)})^{-1/(m-1)}}{\sum_{j=1}^{c}({u^{\prime}}_{ij}^{(t)})^{-1/(m-1)}}\quad\forall i\in[n],\forall j\in[c].$ (18)
Assigning labels and obtaining the clusters. The cluster labels from the algorithm are then obtained by the defuzzification approach of maximum membership value in the finally updated weight matrix $\mathbf{W}$, assigning the vector containing the labels obtained from the clusters as $\mathbf{z}_{n\times 1}$, which is defined as follows,
$z_{i}=\arg\max_{j}(w_{ij})\quad\forall i\in[n],$
therefore, the clusters are $\{Y_{r}\}$, which are defined as $Y_{r}:=\{\mathbf{x}_{i}^{\prime}\mid z_{i}=r\}\quad\forall r\in[c]$.
5 complexity analysis.
In this section, we are going to present the step-by-step complexity analysis of HypeFCM.
Embedding into a Poincaré Disc. Embedding the $n$ datapoints into a Poincaré Disc requires the time complexity of $\mathcal{O}(np)$, where $p$ is the dimension of the points.
Initializing the Weight Matrix. Initializing each row of $\mathbf{W}$ from the Dirichlet distribution requires the time complexity of $\mathcal{O}(c)$. For $n$ rows, the time complexity is $\mathcal{O}(nc)$.
Computing the Centroids. For each of the $c$ centroids, computing $(w_{ij})^{m}$ for all data points requires $\mathcal{O}(n)$. Each centroid update involves weighted Fréchet mean computation using the Riemannian logarithm and exponential maps, which requires the time complexity of $\mathcal{O}(p)$ in each case. Thus, the total time complexity for this step becomes $\mathcal{O}(cnp)$.
Figure 1: The t-SNE visualization of real datasets (a) Glass, (b) Ecoli, and (c) Wine respectively, for FCM and HypeFCM clustering methods.
Computing the distance matrix and applying filtration. Computing the hyperbolic distance for $n$ points from the $c$ centroids requires $\mathcal{O}(cn)$ many computations. For each of the $n$ rows, filtering the $k$ nearest elements from the respective centroid requires the total time complexity $\mathcal{O}(nc\log k)$.
Updating the Weight matrix and centroids. We update the weight matrix $\mathbf{W}$ using Equation 18. Computing the reciprocals of the entries of $\mathbf{U^{\prime}}$ and normalizing by the row sums of $\mathbf{U^{\prime}}$ requires the time complexity $\mathcal{O}(nc)$.
Iteration Step. Performing steps $3$-$6$ for each iteration in Algorithm 1 requires time complexity of $\mathcal{O}(ncp+nc\log k)$. Thus, the total time complexity for $T$ iterations becomes $\mathcal{O}(T(ncp+nc\log k))$.
Assigning the labels and obtaining the clusters. For each of the $n$ rows, finding the maximum weight among the $c$ values requires the time complexity of $\mathcal{O}(nc)$.
Therefore, the overall time complexity for our HypeFCM algorithm is $\mathcal{O}(T(ncp+nc\log k))$, equivalent to $\mathcal{O}(Tncp)$ since $k\ll n$, which is comparable to the overall time complexities of the FCM and PCM algorithms.
Table I: Comparison of Clustering Performance across multiple methods, $K$-means, FCM, PCM, P_SFCM, MiniBatchKmeans, HSFC, with our proposed HypeFCM ON $6$ Synthetic datasets, Presented as Means with Standard Deviations.
Datasets Metric FCM PCM P_SFCM $K$-means MiniBatchKmeans HSFC HypeFCM (Ours)
Cure-t1-2000n-2D ARI 0.471 $\pm$ 0.014 0.079 $\pm$ 0.065 0.486 $\pm$ 0.042 0.491 $\pm$ 0.021 0.505 $\pm$ 0.019 0.493 $\pm$ 0.022 0.723 $\pm$ 0.030
NMI 0.574 $\pm$ 0.012 0.074 $\pm$ 0.059 0.640 $\pm$ 0.038 0.747 $\pm$ 0.018 0.774 $\pm$ 0.016 0.762 $\pm$ 0.019 0.781 $\pm$ 0.039
Cure-t2-4k ARI 0.428 $\pm$ 0.018 0.068 $\pm$ 0.040 0.447 $\pm$ 0.056 0.438 $\pm$ 0.035 0.532 $\pm$ 0.023 0.503 $\pm$ 0.037 0.581 $\pm$ 0.007
NMI 0.607 $\pm$ 0.017 0.072 $\pm$ 0.044 0.608 $\pm$ 0.052 0.666 $\pm$ 0.031 0.649 $\pm$ 0.020 0.617 $\pm$ 0.033 0.676 $\pm$ 0.006
Donutcurves ARI 0.358 $\pm$ 0.021 0.070 $\pm$ 0.035 0.498 $\pm$ 0.037 0.420 $\pm$ 0.032 0.423 $\pm$ 0.031 0.461 $\pm$ 0.019 0.501 $\pm$ 0.006
NMI 0.514 $\pm$ 0.019 0.075 $\pm$ 0.044 0.663 $\pm$ 0.043 0.576 $\pm$ 0.028 0.581 $\pm$ 0.027 0.634 $\pm$ 0.025 0.667 $\pm$ 0.005
Smile1 ARI 0.542 $\pm$ 0.015 0.142 $\pm$ 0.064 0.551 $\pm$ 0.041 0.527 $\pm$ 0.026 0.551 $\pm$ 0.024 0.548 $\pm$ 0.027 0.594 $\pm$ 0.006
NMI 0.606 $\pm$ 0.016 0.325 $\pm$ 0.059 0.612 $\pm$ 0.037 0.559 $\pm$ 0.022 0.607 $\pm$ 0.021 0.607 $\pm$ 0.023 0.701 $\pm$ 0.005
Disk-4000n ARI 0.001 $\pm$ 0.002 0.016 $\pm$ 0.007 0.001 $\pm$ 0.019 -0.002 $\pm$ 0.018 -0.001 $\pm$ 0.025 0.010 $\pm$ 0.025 0.239 $\pm$ 0.033
NMI 0.002 $\pm$ 0.004 0.017 $\pm$ 0.005 0.002 $\pm$ 0.015 0.001 $\pm$ 0.012 0.002 $\pm$ 0.022 0.101 $\pm$ 0.021 0.293 $\pm$ 0.043
3MC ARI 0.584 $\pm$ 0.013 0.142 $\pm$ 0.068 0.553 $\pm$ 0.035 0.508 $\pm$ 0.019 0.546 $\pm$ 0.018 0.498 $\pm$ 0.020 0.586 $\pm$ 0.008
NMI 0.650 $\pm$ 0.012 0.325 $\pm$ 0.062 0.612 $\pm$ 0.033 0.608 $\pm$ 0.017 0.628 $\pm$ 0.016 0.531 $\pm$ 0.018 0.653 $\pm$ 0.009
    Average Rank 4.75 6.33 3.83 4.17 3.25 3.67 1.08
  • indicates a statistically significant difference between the performances of the corresponding algorithm and HypeFCM.
  • indicates the difference between performances of the corresponding algorithm and HypeFCM is not statistically significant.
6 Convergence Analysis.
We will introduce the objective function of HypeFCM. Let $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots\mathbf{x}_{n}\}$ be a sample of $n$ observations in $\mathbb{D}_{\alpha}^{p}$. $\mathbf{V}$ is an $(c\times p)$ matrix that represents the centroids of the $c$ clusters in $\mathbb{D}_{\alpha}^{p}$ and $\mathbf{W}$ is an $(n\times c)$ membership matrix with elements $w_{ij}\in[0,1]$ such that the following least squares criterion is minimized. In this section, we offer insights into the convergence analysis of HypeFCM.
$Minimize\hskip 5.0ptJ_{m}(\mathbf{X},\mathbf{W},\mathbf{V})=\sum_{i=1}^{n}\sum_{j=1}^{c}(w_{ij})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}),$
subject to:
$\sum_{j=1}^{c}w_{ij}=1,\forall i\in[n],$
$w_{ij}\geq 0,\forall i\in[n],j\in[c],$
$\mathbf{x}_{i},\mathbf{v}_{j}\in\mathbb{D}_{\alpha}^{p},\forall i\in[n],j\in[c].$ (19)
6.1 Membership Update
The membership weights $w_{ij}$ are updated as:
$w_{ij}=\frac{d_{H}(\mathbf{x}_{i},\mathbf{v}_{j})^{-2/(m-1)}}{\sum_{l=1}^{c}d_{H}(\mathbf{x}_{i},\mathbf{v}_{l})^{-2/(m-1)}}.$ (20)
6.2 Centroid Update
The hyperbolic centroid update is computed using logarithmic and exponential maps on the Poincaré ball:
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij})^{m}\cdot\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}_{i})}{\sum_{i=1}^{n}(w_{ij})^{m}}\right),$ (21)
where, $\log_{\mathbf{v}}(\cdot)$ and $\exp_{\mathbf{v}}(\cdot)$ are the logarithmic and exponential maps as described in 8, 9.
Lemma 1.
(Bounded distance) For any $\mathbf{x},\mathbf{y}\in\mathbb{D}_{\alpha}^{p}$: $0\leq d_{H}(\mathbf{x},\mathbf{y})<\infty$.
$d_{H}\geq 0$ follows from the properties of $\tanh^{-1}$ and the norm in Equation 7. $d_{H}(x,y)$ is finite, since $\|-\mathbf{x}\oplus_{\alpha}\mathbf{y}\|<\frac{1}{\sqrt{\alpha}}$ in $\mathbb{D}_{\alpha}^{p}$. Detailed proof of this lemma is discussed in the supplementary.
Lemma 2.
For fixed weights $w_{ij}$, there exists a unique minimizer $\mathbf{v}_{j}^{*}$ of the function:
$f(\mathbf{v}_{j})=\sum_{i=1}^{n}(w_{ij})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}).$
Proof of this lemma is given in the supplementary document.
Theorem 1.
Let $J_{m}^{(t)}=J_{m}(\mathbf{W}^{(t)},\mathbf{V}^{(t)})$. The sequence $\{J_{m}^{(t)}\}_{t=0}^{\infty}$ converges to a local minimum.
Proof.
We have shown the convergence analysis of the proposed objective function 19.
Step 1: Descent Property of Membership Update
We aim to show that the HypeFCM algorithm converges to a stationary point of the following objective function defined on the Poincaré ball model $\mathbb{D}_{\alpha}^{p}$: For fixed $\mathbf{V}^{(t)}$, the weight update minimizes $J_{m}$ subject to constraints:
$w_{ij}^{(t+1)}=\frac{d_{H}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}{\sum_{k=1}^{c}d_{H}(\mathbf{x}_{i},\mathbf{v}_{k}^{(t)})^{-2/(m-1)}}.$
For each point $i$, form the Lagrangian from Equation 19.
$L_{i}(w_{i1},...,w_{ic},\lambda_{i})=\sum_{j=1}^{c}(w_{ij})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})+\lambda_{i}(\sum_{j=1}^{c}w_{ij}-1).$ (22)
Taking partial derivatives with respect to $w_{ij}$:
$\frac{\partial{L_{i}}}{\partial{w_{ij}}}=m(w_{ij})^{m-1}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})+\lambda_{i}=0.$
From the above, the following can be deduced,
$m(w_{ij})^{m-1}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})+\lambda_{i}=0$
$\implies w_{ij}=(-\frac{\lambda_{i}}{md_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})})^{1/(m-1)}.$
Applying the sum Constraint $\sum_{j=1}^{c}w_{ij}=1$:
$\sum_{j=1}^{c}(-\frac{\lambda_{i}}{md_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})})^{1/(m-1)}=1.$
Let $\alpha_{i}=(-\frac{\lambda_{i}}{m})^{1/(m-1)}$, then:
$\alpha_{i}\sum_{j=1}^{c}(d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)}))^{-1/(m-1)}=1.$
Therefore,
$\alpha_{i}=\frac{1}{\sum_{j=1}^{c}(d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)}))^{-1/(m-1)}}.$
Hence, as the final update, the following is obtained:
$w_{ij}^{(t+1)}=\frac{d_{H}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}{\sum_{j=1}^{c}d_{H}(\mathbf{x}_{i},\mathbf{v}_{j}^{(t)})^{-2/(m-1)}}.$
Therefore,
$J_{m}(\mathbf{W}^{(t+1)},\mathbf{V}^{(t)})\leq J_{m}(\mathbf{W}^{(t)},\mathbf{V}^{(t)}).$
Step 2: Descent Property of Centroid Update
For fixed weights $\{w_{ij}^{(t+1)}\}$, the centroid update is given by the Riemannian weighted Fréchet mean:
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}_{i})}{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}}\right),$ (23)
where $\log_{\mathbf{v}}(\cdot)$ and $\exp_{\mathbf{v}}(\cdot)$ are the logarithmic and exponential maps at $\mathbf{v}$ in the Poincaré Disc. model. This update minimizes the cost function with respect to $\mathbf{v}_{j}$ under Riemannian geometry, which conforms with the one in Equation 17. For fixed $\mathbf{W}^{(t+1)}$, the centroid update minimizes
$J_{m}:\mathbf{v}_{j}^{(t+1)}=\operatorname*{argmin}_{\mathbf{v}_{j}\in\mathbb{D}_{\alpha}^{p}}\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}).$
$\because J_{m}(\mathbf{W}^{(t+1)},\mathbf{V}^{(t+1)})\leq J_{m}(\mathbf{W}^{(t+1)},\mathbf{V}^{(t)}).$
Therefore,
$J_{m}^{(t)}\geq J_{m}(\mathbf{W}^{(t+1)},\mathbf{V}^{(t)})\geq J_{m}(\mathbf{W}^{(t+1)},\mathbf{V}^{(t+1)})=J_{m}^{(t+1)}.$ (24)
Step 3: Monotonicity and Convergence
$\{J_{m}^{(t)}\}$ is monotonically decreasing sequence in t by 24 and $J_{m}^{(t)}\geq 0$ by Lemma 1. Therefore, $\{J_{m}^{(t)}\}$ converges to some $J^{*}\geq 0$, by Monotone Convergence theorem [38].
The filtration step maintains the monotonicity as it only considers $k$ nearest points from each of the centroids at each iteration. Let $\mathbf{U}^{\prime(t)}$ be the filtered distance matrix, then:
$\sum_{i=1}^{n}\sum_{j=1}^{c}u^{\prime(t)}_{ij}\leq\sum_{i=1}^{n}\sum_{j=1}^{c}u^{(t)}_{ij}.$
Therefore, the unfiltered objective acts as a pointwise majorizer to the one with filtration.
Under these conditions, and by properties of alternating minimization on manifolds [39], the sequence of updates converges to a stationary point of the objective function.
7 experiments
7.1 Details of the Datasets.
We validate the efficacy of HypeFCM on some synthetic and real-world datasets. Iris, Glass, Ecoli, Wine, Phishing URL, Abalone, Zoo, ORHD (Optical Recognition of Handwritten Digits) datasets are taken from UCI machine learning repository [40]; Flights, MNIST datasets are taken from Kaggle, and cure-t1-2000n-2D, cure-t2-4k, donutcurves, disk-4000n, smile1, 3MC are taken from Clustering Benchmark datasets available at https://github.com/deric/clustering-benchmark. Experiments on some of these datasets are shown in the supplementary.
7.2 Experimental Setup & Baselines.
We carried out experiments on HypeFCM on $6$ synthetic datasets along with $7$ real-world datasets. The performance of HypeFCM is measured by involving two well-known performance metrics, namely Adjusted Rand Index (ARI) [41] and Normalized Mutual Information (NMI) [42].
The notable benchmark methods, namely the PCM [13], FCM [15], P_SFCM [26], $k$-means [1], MinibatchKmeans [43], and HSFC [27] clustering methods are considered for comparison with our proposed HypeFCM. A brief description of the datasets used for experimentation is available in Table III.
Figure 2: Grid search plots of $2$ real-world datasets (a) Ecoli, (b) Wine and $2$ synthetic datasets (c) Smile1, (d) Cure-t1-2000n, for two ranges of curvatures ($\alpha$), 00-$1$ and $1$-$1000$ and different filtration ($k$) values against the performance metric, NMI are presented above.
7.3 Parameter Setting.
The performance of HypeFCM depends on three key parameters: the curvature parameter ($\alpha$) of the Poincaré Disc, the value of the filtration ($k$), and the fuzziness parameter ($m$).
We have performed two separate $3$D grid search experiments for each dataset with two parameters (curvature ($\alpha$), filtration value ($k$)) against the performance metric, NMI with one varying the curvature ($\alpha$) from 00 to $1$ by increasing the step size of $0.1$ and the other from $1$ to $1000$ by increasing the step size of $100$ in each iteration on two real-world datasets and two synthetic datasets. The grid search plots in Figure 2 show the performance of HypeFCM on two real-world datasets; Ecoli, Wine, and two synthetic datasets; smile1, cure-t1-2000n-2D respectively.
We also performed two separate $2$D plots for NMI with varying curvature ($\alpha$), for the values of filtration ($k$) varying from $1$ to $15$. We plotted them in two separate plots for $k=1,5,10,15$ in figure 3 & 4 for two real datasets: Ecoli, Wine; and two synthetic datasets: smile1, cure-t1-2000n-2D respectively. We analyze the performance from the figures and conclude that not much variation in performance is observed with the choice of curvature ($\alpha$) regarding the best choice of filtration value ($k$).
In all our experiments, we have considered the value of the fuzziness parameter ($m$) as $2$.
Figure 3: The performance metric NMI vs curvature($\alpha$) plots of HypeFCM for two real-world datasets (a) Ecoli, (b) Wine respectively.
7.4 Experiment on Synthetic Datasets.
We have applied HypeFCM on $6$ synthetic datasets. Performances are presented in Table I. Our HypeFCM algorithm underscores almost all contenders in terms of ARI and NMI. The Average Rank in Table I is just the average of the ranks of ARI, and NMI values obtained from the executions of each method for $15$ times, all initialized with the same seeding at a time. We calculate the $p$-values from the Wilcoxon signed-rank test [44] for the methods compared to HypeFCM and represent them in a table in the supplementary document. Smaller $p$-values ($\leq 0.05$) indicate a statistically significant difference between the performances of the corresponding algorithm and HypeFCM, denoted by or else if there is no statistically significant difference, denoted by .
7.5 Experiments on Real Benchmark Datasets.
We applied HypeFCM on $7$ real-world benchmark datasets and the performances are presented in Table II. Beyond the superior performance in synthetic datasets, we also achieved notable performances on real-world datasets as well. We have demonstrated the t-SNE visualization [45] of HypeFCM in Figure 1 for three real-world datasets, Glass, Ecoli, Wine respectively. The Average Rank mentioned in the table is the same as above for all the real-world datasets. denotes a statistically significant difference between the performances of the corresponding algorithm with our proposed HypeFCM, and indicates no significant difference.
Figure 4: The performance metric NMI vs curvature ($\alpha$) plots of HypeFCM (ours) for two synthetic datasets (a) Smile1, (b) Cure-t1-2000n-2D respectively.
Table II: Comparison of Clustering Performance across multiple methods, $K$-means, FCM, PCM, P_SFCM, MiniBatchKmeans, HSFC, with our proposed HypeFCM ON $7$ Real-World datasets, presented as Means with Standard Deviations.
Datasets Metric FCM PCM P_SFCM $K$-means MiniBatchKmeans HSFC HypeFCM (Ours)
Iris ARI 0.810 $\pm$ 0.012 0.468 $\pm$ 0.071 0.680 $\pm$ 0.058 0.742 $\pm$ 0.023 0.800 $\pm$ 0.015 0.742 $\pm$ 0.021 0.812 $\pm$ 0.015
NMI 0.815 $\pm$ 0.015 0.548 $\pm$ 0.065 0.725 $\pm$ 0.051 0.756 $\pm$ 0.027 0.788 $\pm$ 0.019 0.756 $\pm$ 0.025 0.817 $\pm$ 0.014
Glass ARI 0.228 $\pm$ 0.045 0.005 $\pm$ 0.010 0.129 $\pm$ 0.022 0.270 $\pm$ 0.041 0.210 $\pm$ 0.047 0.264 $\pm$ 0.052 0.598 $\pm$ 0.007
NMI 0.334 $\pm$ 0.038 0.035 $\pm$ 0.016 0.156 $\pm$ 0.026 0.428 $\pm$ 0.036 0.356 $\pm$ 0.042 0.398 $\pm$ 0.047 0.609 $\pm$ 0.016
Ecoli ARI 0.378 $\pm$ 0.041 0.390 $\pm$ 0.047 0.293 $\pm$ 0.028 0.384 $\pm$ 0.039 0.390 $\pm$ 0.038 0.424 $\pm$ 0.057 0.538 $\pm$ 0.006
NMI 0.472 $\pm$ 0.034 0.349 $\pm$ 0.061 0.324 $\pm$ 0.034 0.534 $\pm$ 0.035 0.545 $\pm$ 0.031 0.547 $\pm$ 0.049 0.591 $\pm$ 0.010
Wine ARI 0.366 $\pm$ 0.032 0.042 $\pm$ 0.019 0.287 $\pm$ 0.015 0.352 $\pm$ 0.042 0.365 $\pm$ 0.036 0.375 $\pm$ 0.038 0.385 $\pm$ 0.006
NMI 0.425 $\pm$ 0.036 0.089 $\pm$ 0.024 0.320 $\pm$ 0.019 0.423 $\pm$ 0.037 0.430 $\pm$ 0.032 0.428 $\pm$ 0.034 0.486 $\pm$ 0.006
Zoo ARI 0.527 $\pm$ 0.025 0.302 $\pm$ 0.054 0.157 $\pm$ 0.067 0.709 $\pm$ 0.022 0.680 $\pm$ 0.044 0.447 $\pm$ 0.062 0.726 $\pm$ 0.003
NMI 0.543 $\pm$ 0.029 0.457 $\pm$ 0.049 0.376 $\pm$ 0.058 0.782 $\pm$ 0.025 0.747 $\pm$ 0.039 0.482 $\pm$ 0.058 0.788 $\pm$ 0.005
Abalone ARI 0.144 $\pm$ 0.062 $0$$\pm$ 0.002 0.161 $\pm$ 0.019 0.106 $\pm$ 0.069 0.129 $\pm$ 0.059 0.103 $\pm$ 0.033 0.162 $\pm$ 0.010
NMI 0.134 $\pm$ 0.057 0.002 $\pm$ 0.002 0.163 $\pm$ 0.027 0.099 $\pm$ 0.072 0.126 $\pm$ 0.053 0.097 $\pm$ 0.038 0.159 $\pm$ 0.009
Phishing (5k) ARI $0$$\pm$ 0.003 $0$$\pm$ 0.015 0.060 $\pm$ 0.046 0.001 $\pm$ 0.001 0.001 $\pm$ 0.002 0.002 $\pm$ 0.001 0.169 $\pm$ 0.003
NMI 0.010 $\pm$ 0.005 0.003 $\pm$ 0.008 0.069 $\pm$ 0.031 0.001 $\pm$ 0.007 0.003 $\pm$ 0.001 0.005 $\pm$ 0.002 0.207 $\pm$ 0.002
    Average Rank 3.86 6.29 4.54 4.09 3.30 3.64 1.12
  • indicates a statistically significant difference between the performances of the corresponding algorithm and HypeFCM.
  • indicates the difference between performances of the corresponding algorithm and HypeFCM is not statistically significant.
7.6 Ablation Studies.
We conducted an ablation study by varying the key parameter Filtration.
Figure 5: Comparison of Clustering Performance, NMI vs $k$ with and without filtration for two real-world (a) Wine, (b) Ecoli and two synthetic (c) cure-t1-2000n-2D, (d) smile1 datasets respectively.
HypeFCM with and without Filtration. We conducted two separate experiments, one with filtration and the other without using any filtration on two real-world datasets, Wine, Ecoli and two synthetic datasets, cure-t1-2000n and smile1 respectively. We have performed these experiments by setting the curvature parameter $(\alpha)$ as $1$. Figure 5 shows that using filtration improves the performance of HypeFCM compared to the same without filtration.
Table III: Details of the Datasets Used for the Experimentation.
Datasets No. of samples Dimensions No. of classes
PhiUSIIL Phishing URL 235795 54 2
Zoo 101 16 7
Abalone 4177 8 2
Iris 150 4 2
Ecoli 336 7 8
Wine 178 13 3
Glass 214 9 7
Cure-t1-2000n-2D 2000 2 6
Cure-t2-4k 4200 2 7
Donutcurves 1000 2 4
Smile1 1000 2 4
Disk-4000n 4000 2 2
3MC 400 2 3
8 Discussions.
The performance results of HypeFCM demonstrate its effectiveness across both synthetic and real-world datasets. The algorithm’s superior performance in terms of ARI and NMI metrics against competing methods validates the fundamental hypothesis that hyperbolic geometry can better capture complex hierarchical relationships in clustering tasks.
The significant performance of HypeFCM on $7$ real-world benchmark datasets, as visualized through t-SNE for Glass, Ecoli and Wine datasets, demonstrates how effectively HypeFCM preserves cluster boundaries and captures the inherent hierarchical structures that often characterize real-world datasets. Integrating hyperbolic geometry via the Poincaré Disc model represents a significant methodological advancement. The model’s exponentially expanding space toward its boundary naturally accommodates hierarchical structures that are prevalent in many real-world datasets but challenging to represent in Euclidean space. This geometric advantage, combined with fuzzy clustering principles, allows HypeFCM to capture subtle hierarchical relationships that traditional fuzzy clustering methods cannot properly identify.
9 Conclusion & Future works.
Our proposed HypeFCM marks a significant advancement in handling complex hierarchical data structures by seamlessly integrating hyperbolic geometry with the principles of fuzzy clustering. Our method exploits the unique properties of the Poincaré Disc model, which provides exponentially more space toward its boundary, making it naturally suited for representing hierarchical relationships. Incorporating a weight-based filtering mechanism enhances the algorithm’s ability to focus on relevant geometric relationships while pruning less significant connections, thereby improving both computational efficiency and clustering accuracy. The algorithm’s iterative optimization process, combining hyperbolic distance calculations with fuzzy membership updates, demonstrates robust performance in capturing complex structures that traditional Euclidean-based methods struggle to represent adequately. Experimental results validate the effectiveness of our approach, showing superior performance in preserving hierarchical relationships while maintaining computational feasibility. While the current results are promising, expanding this framework to handle streaming data and exploring integration with deep learning architectures may further enhance the algorithm’s applicability across diverse domains where hierarchical relationships are intrinsic to the data structure.